Given a system like the one you used in lab and the following numbers for the moment of inertia of the T, the radius of the vertical bar and the hanging mass calculate the expected acceleration of the system.
I = 0.040 kg m2
r = 0.015 m
m = 0.500 kg
1
Expert's answer
2015-01-21T10:51:33-0500
The forces on mass m are tension andgravity, so ussing Newton’s second law and expression of moment of inertia we
have:
I =( mgr2/α) − mr2
So accelerationof the system α=mgr2/(I+mr2)=0.03m/s2